# Real-Time Dynamics in Two Dimensions with Tensor Network States via Time-Dependent Variational Monte Carlo

### v3

@@ III.3 Small-$o$ trick in minSR
In tVMC, the memory bottleneck is storing $O$, which can be a serious issue on GPUs. For a PEPS with local dimension $d$, this “small-oo” trick reduces memory cost by a factor of $d$. For the $\mathbb{Z}_{N}$ gauge-invariant PEPS used in Sec. IV.3, the reduction is very substantial—a factor of $dN^{3}$. In practice, this saving often determines whether a large-scale GPU simulation is feasible at all.
@@ III.4 Comparison with double-layer methods
Unlike MPS in 1D [1], PEPS does not possess a canonical form that fixes the gauge freedoms. Ref. [23] introduces local gauge fixing for full-update gate applications, but no globally consistent gauge-fixing procedure is known. In fact, as an alternative, a variational subfamily — isometric PEPS (isoPEPS) [54, 43, 49] — is introduced to explicitly preserve the isometric gauge at the expense of representative expressiveness. However, the necessary shifting of isometric centers introduces significant errors and leads to suboptimal results. In contrast, PEPS–tVMC achieves global gauge fixing efficiently through a single QR projection or simply using the minSR formulation, without any loss of representational power.
@@
In the end, we mention another reason why tVMC may be easier than double-layer methods. Any conservation law in the computational basis can be implemented, in VMC, by sampling in the subsector of configurations with a definite total charge. This is equivalent to acting on the state with a global projector of the desired charge sector, which significantly increases the representability of the underlying bare PEPS.
@@ IV Results
We now apply PEPS–tVMC to study four distinct two-dimensional models—bosonic and fermionic, free and interacting, and including systems with gauge symmetry or nontrivial topology. PEPS configurations are sampled using the sequential sampling algorithm, which was first introduced in Ref. [29], and later proved to satisfy detailed balance in Ref. [55]. The local energies and log-derivatives are computed via the contraction of the single-layer network shown in Fig. 1(d) (see Refs. [29, 55, 28] for methodological details). For Hamiltonians with no explicit time-dependence, the energies are observed to be well-conserved, with relative drift of $2\times 10^{-4}$ to $2\times 10^{-3}$ for the entire trajectory. We show the energy evolution and the simulation parameters in the Supplementary Material [53].
@@ IV.2 Bosonic fractional quantum Hall effect
This result provides a direct real-time confirmation of the composite-fermion prediction of fractional quasiparticle charge. While previous numerical studies of fractional Chern insulators relied primarily on static diagnostics, such as entanglement spectra or density profiles of pinned quasiholes, PEPS–tVMC enables us to directly visualize the transport of fractional charge in two dimensions. This establishes a dynamical counterpart to the usual static signatures of topological order. It would be exciting to study other anyonic dynamics with PEPS–tVMC in the future.
@@ IV.4 Superfluid bosons
IV.5 Error in solving the SR equation
An important factor underlying the accuracy and efficiency of PEPS–tVMC is the numerical stability with which the SR equation—or its reduced minSR variant—can be solved. All simulations in this work solve the SR equation using a Cholesky decomposition [75] of the SR matrix with a diagonal regulator of magnitude $10^{-8}$. Compared to the rank-revealing singular value decomposition (SVD) solver, the Cholesky approach is substantially faster and more parallelizable on GPUs, but it lacks the robustness of SVD when applied to ill-conditioned matrices. Recent NQS–tVMC studies often resort to the SVD solver for this reason [14, 17, 19], trading computational efficiency for numerical robustness. The fact that Cholesky remains stable in all simulations indicates that once gauge redundancies are analytically removed, the PEPS variational manifold becomes exceptionally stable for real-time local quench dynamics.
@@ IV.5 Error in solving the SR equation
To quantify the accuracy of the Cholesky solution, we monitor the SR residual with the regulator removed,
@@
which measures the relative error in satisfying the SR equation. As shown in Fig. 6, the residual remains exceptionally small—typically between $10^{-9}$ and $10^{-25}$, depending on the model. These results confirm the effectiveness of the preconditioning strategies introduced earlier, ensuring that the SR step does not contribute appreciable error during real-time evolution.
@@ V Discussion and Outlook
In this work we have shown that projected entangled pair states, when combined with time-dependent variational Monte Carlo, provide a stable and efficient framework for simulating real-time dynamics in two-dimensional quantum many-body systems. The key technical insight is that the ill-conditioning traditionally associated with PEPS—stemming from gauge redundancy—can be removed analytically in closed form. Once gauge directions are eliminated, the SR equation can be solved efficiently with high numerical stability over long evolution times. This stability, together with the locality and single-layer contraction structure of PEPS within VMC, makes PEPS–tVMC a powerful classical method for studying low-energy dynamics in 2D.
@@ Acknowledgements.
Part of the algorithm is tested using the TenPy code base [77, 78]. The code for production is implemented with JAX [79]. I am supported by a startup grant from the IOP-CAS. I thank my colleagues Yu-Xiang Zhang, Shi-Xin Zhang, and Tao Xiang for sharing computational resources with me. I benefited greatly from discussion and collaboration on various related projects with Bartholomew Andrews, Zhehao Dai, Jannes Nys, Sheng-Hsuan Lin, Sajant Anand, Taige Wang, Michael Zaletel, Frank Pollmann, Wen-Yuan Liu, Jiahang Hu, Jiajun Yu, and Yuntian Gu. After the paper appeared as a preprint, I benefited from questions, comments, and discussion from Giuseppe Carleo, Markus Schmitt, Miles Stoudenmire, Joseph Tindall, Jad Halimeh, Qi Yang, and Yijia Wang. The data that support the findings of this article are openly available [80]; embargo periods may apply.
@@
One needs to form the minSR matrix $T=\mathcal{O}\mathcal{O}^{\dagger}$:
@@ 5 OOdag : (Ns, Ns), obtained from get_OOdag()
Here we document the simulation details of the tVMC carried out in the main text. The integrator of all simulations is 4th-order Runge-Kutta. The regulator in solving the SR or the minSR equation using the Cholesky decomposition is all $10^{-8}$.
@@
Model $D$ $D^{\prime}$ $\mathrm{d}t$ $T$ $N_{s}$ $N_{p}$ $t_{\rm wall}$ $T_{\rm wall}$ SR or minSR
$12\times 12$ Chern insulator 4 16 0.01 12 40960 28224 399s 5.5 days SR
@@
We use energy conservation to check the correctness of our simulations.

### v4

@@ III.3 Small-$o$ trick in minSR
In tVMC, the memory bottleneck is storing $O$ or $S$, which can be a serious issue on GPUs. For a PEPS with local Hilbert space dimension $d$, this “small-oo” trick reduces memory cost by a factor of $d$. For the $\mathbb{Z}_{N}$ gauge-invariant PEPS used in Sec. IV.3, the reduction is very substantial—a factor of $dN^{3}$. In practice, this saving often determines whether a large-scale GPU simulation is feasible at all.
@@ III.4 Comparison with double-layer methods
Unlike MPS in 1D [1], PEPS does not possess a canonical form that fixes the gauge freedoms. Ref. [26] introduces local gauge fixing for full-update gate applications, but no globally consistent gauge-fixing procedure is known. In fact, as an alternative, a variational subfamily — isometric PEPS (isoPEPS) [57, 46, 52] — is introduced to explicitly preserve the isometric gauge at the expense of expressiveness. However, the necessary shifting of isometric centers introduces significant errors and leads to suboptimal results. In contrast, PEPS–tVMC achieves global gauge fixing efficiently through a single QR projection or simply using the minSR formulation, without any loss of representational power.
@@
In the end, we mention another reason why tVMC may be easier than double-layer methods. In VMC, any conservation law that is diagonal in the computational basis can be trivially implemented by sampling configurations in the targeted subsector, with a definite total charge. This is equivalent to acting on the state with a global projector of the desired charge sector, which significantly increases the representability of the underlying bare PEPS.
@@ IV Results
We now apply PEPS–tVMC to study the local quench dynamics of four distinct two-dimensional models—bosonic and fermionic, free and interacting, and including systems with gauge symmetry or nontrivial topology. PEPS configurations are sampled using the sequential sampling algorithm, which was first introduced in Ref. [32], and later proved to satisfy detailed balance in Ref. [58]. The local energies and log-derivatives are computed via the contraction of the single-layer network shown in Fig. 1(d) (see Refs. [32, 58, 31] for methodological details). For Hamiltonians with no explicit time-dependence, the energies are observed to be well-conserved, with relative drift of $2\times 10^{-4}$ to $2\times 10^{-3}$ for the entire trajectory. We show the energy evolution and the simulation parameters in the Supplementary Material [56].
@@ IV.2 Bosonic fractional quantum Hall effect
This result provides a direct real-time confirmation of the composite-fermion prediction of fractional quasiparticle charge. While previous numerical studies of fractional Chern insulators relied primarily on static diagnostics, such as entanglement spectra or density profiles of pinned quasiholes, PEPS–tVMC enables us to directly visualize the transport of fractional charge in two dimensions. This establishes a dynamical counterpart to the usual static signatures of topological order. It would be exciting to study anyonic dynamics in other settings with PEPS–tVMC in the future.
@@ IV.4 Superfluid bosons
IV.5 Error in solving the TDVP equation
An important factor underlying the accuracy and efficiency of PEPS–tVMC is the numerical stability with which the TDVP equation—or its reduced minSR variant—can be solved. All simulations in this work solve the TDVP equation using a Cholesky decomposition [78] of the QGT matrix with a Tikhonov regulator of magnitude $10^{-8}$. Compared to the rank-revealing singular value decomposition (SVD) solver, the Cholesky approach is substantially faster and more parallelizable on GPUs, but it lacks the robustness of SVD when applied to ill-conditioned matrices. Recent NQS–tVMC studies often resort to the SVD solver for this reason [13, 15, 22], trading computational efficiency for numerical robustness. The fact that Cholesky remains stable in all simulations indicates that once gauge redundancies are analytically removed, the PEPS variational manifold becomes exceptionally stable for real-time local quench dynamics.
@@ IV.5 Error in solving the TDVP equation
To quantify the accuracy of the Cholesky solution, we monitor the tVMC residual with the regulator removed,
@@
which measures the relative error in satisfying the TDVP equation. As shown in Fig. 6, the residual remains exceptionally small—typically between $10^{-9}$ and $10^{-25}$, depending on the model. These results confirm the effectiveness of our robust strategies to solve the linear TDVP equations, ensuring that the tVMC step does not contribute appreciable error during real-time evolution.
@@ V Discussion and Outlook
In this work we have shown that projected entangled pair states, when combined with time-dependent variational Monte Carlo, provide a stable and efficient framework for simulating real-time low-energy dynamics in two-dimensional quantum many-body systems. The key technical insight is that the ill-conditioning traditionally associated with PEPS—stemming from gauge redundancy—can be removed analytically in closed form. Once gauge directions are eliminated, the TDVP equation can be solved efficiently with high numerical stability over long evolution times. This stability, together with the locality and single-layer contraction structure of PEPS within VMC, makes PEPS–tVMC a powerful classical method for studying low-energy dynamics in 2D.
@@ Acknowledgements.
Part of the algorithm is tested using the TenPy code base [80, 81]. The code for production is implemented with JAX [82]. Y.W. is supported by a startup grant from the IOP-CAS. Y.W. thanks his colleagues Yu-Xiang Zhang, Shi-Xin Zhang, and Tao Xiang for sharing computational resources with him. He benefited greatly from discussion and collaboration on various related projects with Bartholomew Andrews, Zhehao Dai, Sheng-Hsuan Lin, Sajant Anand, Taige Wang, Michael Zaletel, Frank Pollmann, Wen-Yuan Liu, Jiahang Hu, Jiajun Yu, and Yuntian Gu. J.N. thanks Juan Carrasquilla for insightful discussions. After the paper appeared as a preprint, we benefited from questions, comments, and discussion from Giuseppe Carleo, Markus Schmitt, Miles Stoudenmire, Joseph Tindall, Jad Halimeh, Qi Yang, and Yijia Wang. The data that support the findings of this article are openly available [83]; embargo periods may apply. The simulations were supported by the IOP-CAS computing facilities.
@@
One needs to form the minSR matrix (also referred to as the neural tangent kernel (NTK)), $T=\mathcal{O}\mathcal{O}^{\dagger}$:
@@ 5 OOdag : (Ns, Ns), obtained from get_OOdag()
Here we document the simulation details of the tVMC carried out in the main text. The integrator of all simulations is 4th-order Runge-Kutta. The Tikhonov regulator in solving the tVMC or the minSR equation using the Cholesky decomposition is $10^{-8}$ everywhere.
@@
Model $D$ $D^{\prime}$ $\mathrm{d}t$ $T$ $N_{s}$ $N_{p}$ $t_{\rm wall}$ $T_{\rm wall}$ TDVP or minSR
$12\times 12$ Chern insulator 4 16 0.01 12 40960 28224 399s 5.5 days TDVP
@@
We use energy conservation to check the correctness of our simulations in Figs. 1(a), 1(c), and 1(d). It is worth noting that the TDVP equations fulfil energy conservation by construction, even for ill-estimated preconditioners (see e.g. Ref. [13]). Hence, predominant sources of violating energy conservation are the finite step size $dt$, the Runge-Kutta integrator, and a limited sample size.
The energy in the bosonic Hofstadter model is also shown in Fig. 1(b), but is not expected to be conserved due to the pumping. As a consistency check, we show that there is little effect of the bond dimension of the boundary MPS.